# Conformal Predictive Systems Under Covariate Shift

### v1

@@ Abstract
Keywords Conformal prediction  $\cdot$ Conformal predictive systems  $\cdot$ Predictive distributions  $\cdot$ Regression  $\cdot$ Covariate shift
@@ 1 Introduction
Conformal Predictive Systems (CPS) are a relatively recent development in Conformal Prediction (CP) (Vovk et al., 2019, 2020a). CPS construct predictive distributions by arranging p-values into a nonparametric probability distribution. This distribution satisfies a finite-sample property of validity under the Independent and Identically Distributed (IID) model, i.e., the observations are produced independently from the same probability measure. CPS can be seen as a generalization of point and conformal regressors since they can easily produce point predictions and prediction intervals by leveraging the generated predictive distributions. They allow for more informative and trustworthy decision-making (Vovk et al., 2018).
@@
The exchangeability assumption, which allows for provably valid inference for CP and is a weaker assumption than the IID assumption (Shafer and Vovk, 2008), and similarly, the IID assumption for CPS, are standard in machine learning. However, distributional shifts between training and inference data are common in time series, counterfactual inference, and machine learning for scientific discovery but violate these assumptions. While a growing amount of literature has been contributed to extending CP beyond the exchangeability assumptions (Tibshirani et al., 2019; Gibbs and Candes, 2021; Prinster et al., 2022; Yang et al., 2022; Gibbs and Candès, 2023), allowing (conservatively) valid inference under various types of distributional shifts, no contribution has been made towards extending CPS beyond the IID model. For example, in treatment effect estimation, this extension could allow calibrated predictive distribution beyond the randomized trial setting (Jonkers et al., 2024b), as in a nonrandomized setting, the covariate distributions for treated and control subjects differ from the target population. Therefore, this work extends CPS beyond the IID model by proposing weighted CPS that constructs valid nonparametric predictive distributions for problems where the covariate distributions of the training and testing data differ, assuming their likelihood ratio is known or can be estimated.
@@ 2.3 Weighted Conformal Prediction
Tibshirani et al. (2019) was one of the first works to extend conformal prediction beyond the exchangeability assumption to deal with covariate shifts. Specifically, they propose Weighted Conformal Prediction (WCP) to deal with covariate shifts where the likelihood ratio between the training $P_{X}$ and test $\tilde{P}_{X}$ covariate distributions is known. In WCP, the empirical distribution of nonconformity scores at the training points gets reweighted, and thus each nonconformity score $R_{i}$ gets weighted by a probability $p_{i}^{w}(x)$ proportional to the likelihood ratio $w(x_{i})=\frac{d\tilde{P}_{X}(x_{i})}{P_{X}(x_{i})}$:
@@
This results in an adjusted empirical distribution of nonconformity scores depicted in Table 1.
@@
Regular Weighted
$\frac{1}{n+1}\sum_{i=1}^{n}\delta_{R_{i}}+\frac{1}{n+1}\delta_{\infty}$ $\sum_{i=1}^{n}p_{i}^{w}(x)\delta_{R_{i}}+p_{n+1}^{w}(x)\delta_{\infty}$
@@
where $w(x_{1:n})=(w(x_{1}),...,w(x_{n}))$. Note that it is possible to learn the likelihood ratio $w(x_{i})=\frac{d\tilde{P}_{X}(x_{i})}{P_{X}(x_{i})}$ between training and test covariate distribution, as showed by Tibshirani et al. (2019), if it is reasonably accurate.
@@ 2.4 Conformal Predictive Systems
Conformal Predictive Systems (CPS) allow the construction of predictive distributions by extending upon full CP. CPS produces conformal predictive distributions by arranging p-values into a probability distribution function (Vovk et al., 2019). These p-values are created with the help of specific types of conformity measures. Vovk et al. (2019) defines a CPS as a function that is both a conformal transducer (Definition 1) and a Randomized Predictive System (RPS) (Definition 2).
Definition 1 (Conformal Transducer, Vovk et al. (2022)).
@@
    For each training sequence $(z_{1},...,z_{n})\in\textbf{Z}^{n}$ and test object $x\in\textbf{X}$, the function $Q(z_{1},...,z_{n},(x,y),\tau)$ is monotonically increasing both in $y$ and $\tau$. Put differently, for each $\tau\in[0,1]$, the function
    $y\in\mathbb{R}\rightarrow Q(z_{1},...,z_{n},(x,y),\tau)$
@@
    For each $\tau,\tau^{\prime}\in[0,1]$ and each test object $x\in\textbf{X}$,
    $Q(z_{1},...,z_{n},(x,y),\tau)>Q(z_{1},...,z_{n},(x,y^{\prime}),\tau^{\prime}),\qquad\text{if}\quad y>y^{\prime}$
@@
    For each training sequence $(z_{1},...,z_{n})\in\textbf{Z}^{n}$ and test object $x\in\textbf{X}$,
    $\lim_{y\rightarrow-\infty}Q(z_{1},...,z_{n},(x,y),0)=0$
@@
    As a function of random training observations $z_{1}\sim P,...,z_{n}\sim P$, and a random number $\tau\sim Uniform(0,1)$, all assumed to be independent, the distribution of Q is uniform:
@@
Definition 2 that defines an RPS is in verbatim from Vovk et al. (2019), except requirement R1.2, which is appended to the definition as we believe this is a requirement which is implicitly assumed by Vovk et al. (2019).
@@ 3 Weighted Conformal Predictive System
As WCP extends upon CP, we propose to reweigh the conformity scores with a probability $p_{i}^{w}(x)$ proportional to the likelihood ratio $w(x_{i})=\frac{d\tilde{P}_{X}(x_{i})}{P_{X}(x_{i})}$, to present a weighted conformal transducer where the output is defined by conformity measure $A$ and likelihood ratio $w(x)=\frac{d\tilde{P}_{X}(x)}{P_{X}(x)}$,
$Q(z_{1},...,z_{n},\frac{d\tilde{P}}{P},(x,y),\tau):=\sum_{i=1}^{n+1}[R_{i}^{y}<R_{n+1}^{y}]p_{i}^{w}(x)+\sum_{i=1}^{n+1}[R_{i}^{y}=R_{n+1}^{y}]p_{i}^{w}(x)\tau$ (11)
@@
A function is a Weighted Conformal Predictive System (WCPS) if it is both a weighted conformal transducer and an RPS. To prove that under specific conformity measures $A$, e.g., monotonic conformity measures, a weighted conformal transducer is also an RPS, we need to prove Conjecture 1, i.e., that the weighted conformal transducer is probabilistically calibrated.
Conjecture 1.
@@
$\forall\alpha\in[0,1]:\mathbb{P}_{z_{1:n}\sim P,z_{n+1}\sim\tilde{P}}\{Q(z_{1},...,z_{n},\frac{d\tilde{P}}{P},z_{n+1},\tau)\leq\alpha\}=\alpha$ (12)
We leave this proof for future work. However, if proven, Conjecture 2 can be easily proven by following the same procedure as the proof of Lemma 1 in Vovk (2022) using Conjecture 1 instead of the property of validity of a conformal transducer.
Conjecture 2 (Weighted Version of Lemma 1 in Vovk (2022)).
@@
Then, the weighted conformal transducer corresponding to $A$ is an RPS.
In other words, a weighted conformal transducer based on a monotonic conformity measure satisfying the aforementioned requirements is also an RPS.
@@ 3.1 Weighted Split Conformal Predictive Systems
$Q(z_{1},...,z_{n},\frac{d\tilde{P}}{P},(x,y),\tau):=\sum_{i=m+1}^{n}[R_{i}<R^{y}]p_{i}^{w}(x)+\sum_{i=m+1}^{n}[R_{i}=R^{y}]p_{i}^{w}(x)\tau+p^{w}_{n+1}(x)\tau$ (15)
Similarly to WCPS, a function is a Weighted Split Conformal Predictive System (WSCPS) if it is both a split conformal transducer and a randomized predictive system. Thus, we also need to prove a notion of validity in the form of calibration in probability, see Conjecture 3. We leave this proof for future work, but we show in Section 4 with simulation experiments that this empirically seems to be the case.
Conjecture 3.
@@
$\forall\alpha\in[0,1]:\mathbb{P}_{z_{m+1:n}\sim P,z_{n+1}\sim\tilde{P}}\{Q(z_{1},...,z_{n},\frac{d\tilde{P}}{P},z_{n+1},\tau)\leq\alpha\}=\alpha$ (16)
Conjecture 4.
The weighted split conformal transducer (15) is an RPS if and only if it is based on a balanced isotonic split conformity measure.
A proof of Conjecture 4 will follow the same procedure as the proof of Proposition 1 and 2 in Vovk et al. (2020a) using Conjecture 3 instead of the property of validity of a split conformal transducer.
@@ 4 Experiments
We evaluate (weighted) CPS under a covariate shift on empirical and synthetic data, and use (weighted) split CPS approaches for efficiency. For implementing WSCPS, we made an extension of the python package crepes (Boström, 2022), named weighted-crepes. A more detailed description can be found in Appendix B. The Python code to reproduce the simulation results can be found at https://github.com/predict-idlab/crepes-weighted.
@@ 4.2 Results
Finally, we validate by simulation Conjecture 3 by producing p-values with the (W)SCPS by setting $y$ to the label $y_{n+1}$ and checking if their histogram follows a uniform distribution. In the probabilistic forecasting literature, this is often referred to as Probability Integral Transforms (PIT) histograms (Gneiting et al., 2007).
@@
The results are depicted in Figure 1. We observe similar results as WCP (Tibshirani et al., 2019); in row 1 of Figure 1) we observe undercoverage for SCPS under covariate shift. The WSCPS brings the average coverage to the desired level under covariate shift for both experiments, while the SCPS constructed intervals considerably undercover; see row 2 of Figure 1. We also observe that the heuristic for the reduced (effective) calibration set size due to the weighting operation of WCP, see Equation 8, is also a good heuristic for WSCPS. This is shown in the third row of Figure 1, where we observe similar dispersion of coverage over experiment trials for WSCPS and SCPS with a reduced calibration set.
(a) Airfoil data
(b) Synthetic data
@@
Figure 2 shows the performance of different SCPS in terms of CRPS across the different trials. We see a performance difference when a covariate shift is present and not. The WSCPS consistently (slightly) outperforms the SCPS under covariate shift for both datasets. However, it is difficult to see in the second row of Figure 2. Therefore, we also perform a post-hoc Friedman-Nemenyi test (see Figure 3). The SCPS under no shift with a calibration set size equal to the effective sample size of WSCPS has a significantly better CRPS score than WSCPS. This is expected since under covariate shift, the model $\hat{\mu}$ is trained on training data differently distributed as the test set, as Tibshirani et al. (2019) also indicated. Ideally, $\hat{\mu}$ should be adjusted for the covariate shift; however, we leave this for future work.
(a) Airfoil data
(b) Synthetic data
@@
We validate by simulation Conjecture 3, which states that under covariate shift, the weighted split conformal transducer produced $p$-values are distributed uniformly on $[0,1]$ when we know the likelihood ratio of the covariate distribution of the training and test set. The results of the simulation experiments, depicted in Figure 4, indicate that Conjecture 3 is empirically valid and that it breaks when we do not account for the covariate shift.
@@ 5 Conclusion
In future work, we aim to provide rigorous proofs for the conjectures presented in this paper to establish the theoretical underpinnings of our proposed methods. Additionally, we will evaluate our proposed framework for counterfactual inference and incorporate it into our recently proposed Conformal Monte-Carlo meta-learners (Jonkers et al., 2024b), which opens the possibility of giving validity guarantees for predictive distributions of individual treatment effect beyond the randomized trial setting. Overall, our contributions offer a promising avenue for addressing covariate shifts in predictive modeling, with potential applications in diverse fields such as healthcare, finance, and climate science.
Acknowledgements
Part of this research was supported through the Flemish Government (Flanders AI Research Program).

### v2

@@ Abstract
keywords:
Conformal prediction, Conformal predictive systems, Predictive distributions, Regression, Covariate shift
@@ 1 Introduction
Conformal predictive systems (CPS) are a relatively recent development in Conformal prediction (CP) (Vovk et al., 2019, 2020a). CPS construct predictive distributions by arranging p-values into a nonparametric probability distribution. This distribution satisfies a finite-sample property of validity under the independent and identically distributed (IID) model, i.e., the observations are produced independently from the same probability measure. CPS can be seen as a generalization of point and conformal regressors since they can easily produce point predictions and prediction intervals by leveraging the generated predictive distributions. They allow for more informative and trustworthy decision-making (Vovk et al., 2018). For example, an obvious use case for predictive distributions is to obtain threshold values such that a decision-maker can determine the likelihood of a certain outcome falling within a specific range. This is particularly useful in applications where understanding the probability distribution of predictions is crucial, such as in risk management, medical diagnostics, and financial forecasting. By setting appropriate thresholds based on the predictive distributions, one can make more informed decisions that balance the trade-offs between different types of errors or risks.
@@
The exchangeability assumption, which allows for provably valid inference for CP and is a weaker assumption than the IID assumption (Shafer and Vovk, 2008), and similarly, the IID assumption for CPS, are standard in machine learning. However, distributional shifts between training and inference data are common in time series, counterfactual inference, and machine learning for scientific discovery but violate these assumptions. While a growing amount of literature has been contributed to extending CP beyond the exchangeability assumptions (Tibshirani et al., 2019; Gibbs and Candes, 2021; Prinster et al., 2022; Yang et al., 2022; Gibbs et al., 2023), allowing (conservatively) valid inference under various types of distributional shifts, no contribution has been made towards extending CPS beyond the IID model. For example, in treatment effect estimation, this extension allows for calibrated predictive distribution beyond the randomized trial setting (Jonkers et al., 2024b), as in a nonrandomized setting, the covariate distributions for treated and control subjects differ from the target population. Therefore, this work extends CPS beyond the IID model by proposing weighted CPS (WCPS) that constructs valid nonparametric predictive distributions for problems where the covariate distributions of the training and testing data differ, assuming their likelihood ratio is known or can be estimated.
@@ 2.3 Weighted Conformal Prediction
Tibshirani et al. (2019) was one of the first works to extend conformal prediction beyond the exchangeability assumption to deal with covariate shifts. Specifically, they propose weighted conformal prediction (WCP) to deal with covariate shifts where the likelihood ratio between the training $P_{X}$ and test $\tilde{P}_{X}$ covariate distributions is known. In WCP, the empirical distribution of nonconformity scores at the calibration points gets reweighted, and thus each nonconformity score $R_{i}$ gets weighted by a probability $p_{i}^{w}(x)$ proportional to the likelihood ratio $w(x_{i})=\frac{d\tilde{P}_{X}(x_{i})}{dP_{X}(x_{i})}$:
@@
This results in an adjusted empirical distribution of nonconformity scores depicted in Table LABEL:tab:emp-dist.
[…]
tab:emp-dist Regular Weighted $\frac{1}{n+1}\sum_{i=1}^{n}\delta_{R_{i}}+\frac{1}{n+1}\delta_{\infty}$ $\sum_{i=1}^{n}p_{i}^{w}(x)\delta_{R_{i}}+p_{n+1}^{w}(x)\delta_{\infty}$
@@
where $w(x_{1:n})=(w(x_{1}),...,w(x_{n}))$. Note that it is possible to learn the likelihood ratio $w(x_{i})=\frac{d\tilde{P}_{X}(x_{i})}{dP_{X}(x_{i})}$ between training and test covariate distribution, as showed by Tibshirani et al. (2019), if it is reasonably accurate.
@@ 2.4 Conformal Predictive Systems
Conformal predictive systems (CPS) allow the construction of predictive distributions by extending upon full CP. CPS produces conformal predictive distributions by arranging p-values into a probability distribution function (Vovk et al., 2019). These p-values are created with the help of specific types of conformity measures. Vovk et al. (2019) defines a CPS as a function that is both a conformal transducer (Definition 2.1) and a randomized predictive system (RPS) (Definition 2.2).
Definition 2.1 (Conformal Transducer, Vovk et al. (2022)).
@@
    For each training sequence $(z_{1},...,z_{n})\in\textbf{Z}^{n}$ and test object $x\in\textbf{X}$, the function $Q(z_{1},...,z_{n},(x_{n+1},y),\tau)$ is monotonically increasing both in $y$ and $\tau$. Put differently, for each $\tau\in[0,1]$, the function
    $y\in\mathbb{R}\rightarrow Q(z_{1},...,z_{n},(x_{n+1},y),\tau)$
@@
    For each $\tau,\tau^{\prime}\in[0,1]$ and each test object $x_{n+1}\in\textbf{X}$,
    $Q(z_{1},...,z_{n},(x_{n+1},y),\tau)\geq Q(z_{1},...,z_{n},(x_{n+1},y^{\prime}),\tau^{\prime}),\qquad\text{if}\quad y>y^{\prime}$
@@
    For each training sequence $(z_{1},...,z_{n})\in\textbf{Z}^{n}$ and test object $x_{n+1}\in\textbf{X}$,
    $\lim_{y\rightarrow-\infty}Q(z_{1},...,z_{n},(x_{n+1},y),0)=0$
@@
    As a function of random training observations $z_{1}\sim P,...,z_{n}\sim P$, $z_{n+1}\sim P$, and a random number $\tau\sim Uniform(0,1)$, all assumed to be independent, the distribution of Q is uniform:
@@
Definition 2.2 that defines an RPS is in verbatim from Vovk et al. (2019), except requirement R1.2, which is appended to the definition as we believe this is a requirement which is implicitly assumed by Vovk et al. (2019).
@@ 3 Weighted Conformal Predictive System
As WCP extends upon CP, we propose to reweigh the conformity scores with a probability $p_{i}^{w}(x)$ proportional to the likelihood ratio $w(x_{i})=\frac{d\tilde{P}_{X}(x_{i})}{dP_{X}(x_{i})}$, to present a weighted conformal transducer where the output is defined by conformity measure $A$ and likelihood ratio $w(x)=\frac{d\tilde{P}_{X}(x)}{dP_{X}(x)}$,
$Q(z_{1},...,z_{n},\frac{d\tilde{P}}{dP},(x_{n+1},y),\tau):=\sum_{i=1}^{n+1}[R_{i}^{y}<R_{n+1}^{y}]p_{i}^{w}(x)+\sum_{i=1}^{n+1}[R_{i}^{y}=R_{n+1}^{y}]p_{i}^{w}(x)\tau$ (11)
@@
A function is a weighted conformal predictive system (WCPS) if it is both a weighted conformal transducer and a weighted RPS (WRPS), i.e., an RPS probabilistically calibrated under covariate shift. To prove that under specific conformity measures $A$, e.g., monotonic conformity measures, a weighted conformal transducer is also a WRPS, we need to prove Conjecture 3.1, i.e., that the weighted conformal transducer is probabilistically calibrated.
Conjecture 3.1.
@@
$\forall\alpha\in[0,1]:\mathbb{P}_{z_{1:n}\sim P,z_{n+1}\sim\tilde{P}}\{Q(z_{1},...,z_{n},\frac{d\tilde{P}}{dP},z_{n+1},\tau)\leq\alpha\}=\alpha$ (12)
We leave this proof for future work. However, if proven, Conjecture 3.2 can be easily proven by following the same procedure as the proof of Lemma 1 in Vovk (2022) using Conjecture 3.1 instead of the property of validity of a conformal transducer.
Conjecture 3.2 (Weighted Version of Lemma 1 in Vovk (2022)).
@@
Then, the weighted conformal transducer corresponding to $A$ is a WRPS.
In other words, a weighted conformal transducer based on a monotonic conformity measure satisfying the aforementioned requirements is also a WRPS.
@@ 3.1 Weighted Split Conformal Predictive Systems
$Q(z_{1},...,z_{n},\frac{d\tilde{P}}{dP},(x,y),\tau):=\sum_{i=m+1}^{n}[R_{i}<R^{y}]p_{i}^{w}(x)+\sum_{i=m+1}^{n}[R_{i}=R^{y}]p_{i}^{w}(x)\tau+p^{w}_{n+1}(x)\tau$ (15)
Similarly to WCPS, a function is a weighted split conformal predictive system (WSCPS) if it is both a weighted split conformal transducer and a WRPS. Thus, we also need to prove a notion of validity in the form of calibration in probability, see Conjecture 3.3. We leave this proof for future work, but we show in Section 4 with simulation experiments that this empirically seems to be the case.
Conjecture 3.3.
@@
$\forall\alpha\in[0,1]:\mathbb{P}_{z_{m+1:n}\sim P,z_{n+1}\sim\tilde{P}}\{Q(z_{1},...,z_{n},\frac{d\tilde{P}}{dP},z_{n+1},\tau)\leq\alpha\}=\alpha$ (16)
Conjecture 3.4.
The weighted split conformal transducer (15) is a WRPS if and only if it is based on a balanced isotonic split conformity measure.
A proof of Conjecture 3.4 will follow the same procedure as the proof of Proposition 1 and 2 in Vovk et al. (2020a) using Conjecture 3.3 instead of the property of validity of a split conformal transducer.
@@ 4 Experiments
We evaluate (weighted) CPS under a covariate shift on empirical and synthetic data, and use (weighted) split CPS approaches for efficiency. For implementing WSCPS, we made an extension of the python package crepes (Boström, 2022), named crepes-weighted. A more detailed description can be found in Appendix B. The Python code to reproduce the simulation results can be found at https://github.com/predict-idlab/crepes-weighted.
@@ 4.2 Results
Finally, we validate by simulation Conjecture 3.3 by producing p-values with the (W)SCPS by setting $y$ to the label $y_{n+1}$ and checking if their histogram follows a uniform distribution. In the probabilistic forecasting literature, this is often referred to as probability integral transforms (PIT) histograms (Gneiting et al., 2007).
@@
The results are depicted in Figure LABEL:fig:cov. We observe similar results as WCP (Tibshirani et al., 2019); in row 1 of Figure LABEL:fig:cov) we observe undercoverage for SCPS under covariate shift. The WSCPS brings the average coverage to the desired level under covariate shift for both experiments, while the SCPS constructed intervals considerably undercover; see row 2 of Figure LABEL:fig:cov. We also observe that the heuristic for the reduced (effective) calibration set size due to the weighting operation of WCP, see Equation 8, is also a good heuristic for WSCPS. This is shown in the third row of Figure LABEL:fig:cov, where we observe similar dispersion of coverage over experiment trials for WSCPS and SCPS with a reduced calibration set.
[…]
@@
Figure LABEL:fig:crps-hist shows the performance of different SCPS in terms of CRPS across the different trials. We see a performance difference when a covariate shift is present and not. The WSCPS consistently (slightly) outperforms the SCPS under covariate shift for both datasets. However, it is difficult to see in the second row of Figure LABEL:fig:crps-hist. Therefore, we also perform a post-hoc Friedman-Nemenyi test (see Figure LABEL:fig:crps-tests). The SCPS under no shift with a calibration set size equal to the effective sample size of WSCPS has a significantly better CRPS score than WSCPS. This is expected since under covariate shift, the model $\hat{\mu}$ is trained on training data differently distributed as the test set, as Tibshirani et al. (2019) also indicated. Ideally, $\hat{\mu}$ should be adjusted for the covariate shift; however, we leave this for future work.
[…]
@@
We validate by simulation Conjecture 3.3, which states that under covariate shift, the weighted split conformal transducer produced $p$-values are distributed uniformly on $[0,1]$ when we know the likelihood ratio of the covariate distribution of the training and test set. The results of the simulation experiments, depicted in Figure 4, indicate that Conjecture 3.3 is empirically valid and that it breaks when we do not account for the covariate shift.
@@ 5 Conclusion
In future work, we aim to provide rigorous proofs for the conjectures presented in this paper to establish the theoretical underpinnings of our proposed methods. Additionally, we will evaluate our proposed framework for counterfactual inference and incorporate it into our recently proposed conformal Monte-Carlo meta-learners (Jonkers et al., 2024b), which opens the possibility of giving validity guarantees for predictive distributions of individual treatment effect beyond the randomized trial setting. Moreover, we believe that similar applications and extensions of the WCP (Tibshirani et al., 2019), such as addressing label shift (Podkopaev and Ramdas, 2021), feedback covariate shift (Fannjiang et al., 2022), and survival analysis (Candès et al., 2023), could be applied to WSCPS. Overall, our contributions offer a promising avenue for addressing covariate shifts in predictive modeling, with potential applications in diverse fields such as healthcare, finance, and climate science.
[…]
Part of this research was supported through the Flemish Government (AI Research Program).